# Progressive-Resolution Policy Distillation: Leveraging Coarse-Resolution Simulations for Time-Efficient Fine-Resolution Policy Learning

### v2

@@ I Introduction
We propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD) for achieving time-efficient policy learning of fine-resolution simulations by utilizing coarse-resolution simulations effectively. PRPD progressively improves simulation resolution while learning and transferring policies at each stage. PRPD uses a conservative policy transfer scheme to regularize policies, stabilizing policy transfer across different resolution simulations having different behaviors. We demonstrated the effectiveness of PRPD by constructing a variable-resolution rock excavation simulator using Isaac Gym [21]. This paper demonstrated a 7-fold improvement in learning time efficiency (PRPD: 90 minutes, policy learning in the finest-resolution simulation: 600 minutes), resulting in an 8-hour time gap. This difference is likely to become more significant with increasing state-action space complexity or task difficulty. Furthermore, in evaluations of the learned policies within a real-world environment containing nine types of rocks, PRPD achieved rock excavation in approximately 90% of scenarios.
@@ II-C Sample Efficient Reinforcement Learning Methods
Several methods improve sample efficiency in RL. Model-based RL uses a learned model of the environment’s dynamics [26, 27, 28]. Auxiliary tasks, autoencoders, and contrastive learning extract better features from observations [29, 30, 31]. Data augmentation enhances information content by transforming observation samples [32, 33, 34]. Ensemble learning improves robustness and adaptability with multiple models [31, 35, 36]. Imitation learning trains policies from demonstrations, and reward function learning from small demonstration sets reduces human effort [37, 38, 39]. Off-policy RL such as soft actor-critic (SAC) efficiently explores and learns by reusing past interaction data [40, 41].
@@
Various methods have improved RL’s “sample efficiency” by extracting more information from each sample or increasing update times to maximize sample utilization. However, these approaches often come with computational overhead, such as increased calculation time and resource demands by adding learning components, which can compromise “time efficiency,” the overall training time for policy learning. Since this paper focuses on time efficiency, our experiments utilize on-policy RL, which provides high time efficiency due to its simpler learning structure, fewer model parameters, and reduced update frequency [42, 43].
Our approach can be interpreted as a type of curriculum reinforcement learning (CRL) since it also involves gradually changing the tasks from simple to complex [44, 45]. In detail, CRL gradually increases task difficulty, reducing the sample requirements compared to learning solely on complex tasks. However, to the best of our knowledge, our approach is the first to treat differences in the spatial resolution of complex simulations as distinct tasks.
@@ III-A Reinforcement Learning
$\begin{cases}L_{\theta,\phi}^{\pi,V}\!=\!\!\!&\mathbb{E}_{\mathcal{D}_{i}}[\min\{\rho A_{k}(s,a),\text{clip}(\rho,1-\epsilon,1+\epsilon)A_{k}(s,a)\}\\ &-c_{1}(V_{k}(s)-\hat{R})^{2}+c_{2}\pi_{k}(a|s)\log\pi_{k}(a|s)],\end{cases}$ (3)
@@ III-B Conservative Policy Update in Reinforcement Learning
IV Progressive-Resolution Policy Distillation
Simulator Assumption: To enable policy transfer, our framework follows some simulator assumptions. Policies are designed consistently across different resolutions, even though different simulators typically vary in observations and actions. We assume that resolution-affected elements do not impact observations or actions. The policies’ objectives remain the same, with only the resolution differing, so the reward function is kept identical. This framework supports policy learning in high-resolution simulations, assuming that behavior is less accurate at coarser resolutions and becomes finer as resolution improves.
IV-A Resolution Scheduling
IV-A1 Execution Flow
IV-A2 Scheduling Setups
@@ IV-B2 Policy Update with Conservative Policy Transfer
$\begin{cases}\alpha_{k+1}=\alpha_{0}\ \mathbb{E}_{s\sim\mathcal{D}_{i}^{(n)}}\big{[}\mathbb{E}_{a^{\prime}\sim\pi^{(n-1)}(s)}[Q_{k}^{(n)}(s,a^{\prime})]\\ ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathbb{E}_{a\sim{\pi}_{k}^{(n)}(s)}[Q_{k}^{(n)}(s,{a})]\big{]},\end{cases}$ (7)
where $\alpha_{0}$ is a scaling coefficient, $\mathcal{D}^{(n)}$ is the rollout buffer of the current resolution, and $k$ is an update number. Estimated $\alpha_{k+1}$ is utilized by clipping to $[0,1]$. Finally, the loss function of conservative policy update is defined:
$\begin{cases}&\mathcal{L}_{\theta^{(n)}}^{\pi^{(n)}}=\mathbb{E}_{(s,a)\sim\mathcal{D}_{i}^{(n)}}\big{[}\text{KL}\{\pi^{(n)}_{k}(s,a)||\\ &(1-\alpha_{k+1})\pi^{(n)}_{k}(s,a)+\alpha_{k+1}\pi^{(n-1)}(s,a)\}\big{]},\end{cases}$ (8)
where $\theta^{(n)}$ represents the policy network parameters for the current resolution, and KL denotes the Kullback–Leibler divergence. To estimate $\alpha$, PRPD learns $Q^{(n)}$, the value function parameterized by $\psi^{(n)}$. The value function $Q(s,a)$ is updated using a loss function based on the temporal difference error. The loss function is defined as:
$\begin{cases}\mathcal{L}_{\psi^{(n)}}^{Q^{(n)}}=&\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}_{i}^{(n)}}\big{[}\{Q^{(n)}_{k}(s,a)-\\ &r+\gamma\mathbb{E}_{a^{\prime}\sim\pi^{(n)}(s^{\prime})}[Q^{(n)}_{k}(s^{\prime},a^{\prime})]\}^{2}\big{]}.\end{cases}$ (9)
@@
$\mathcal{L}_{\theta^{(n)},\phi^{(n)},\psi^{(n)}}=L_{\theta^{(n)},\phi^{(n)}}^{\pi^{(n)},V^{(n)}}+c_{3}\mathcal{L}_{\theta^{(n)}}^{\pi^{(n)}}+c_{4}\mathcal{L}_{\psi^{(n)}}^{Q^{(n)}},$ (10)
@@ VI Experiments
We conducted experiments to validate the following objectives. PRPD can learn policies in a shorter learning time compared to previous works (Section VI-B). The conservative policy transfer scheme stabilizes learning in PRPD (Section VI-C). Differences in resolution scheduling affect policy learning (Section VI-D). Trade-off between time efficiency and sample efficiency compared to other RL methods (Section VI-E). The policy trained in the simulator can be applied to various real-world environments (Section VI-F).
@@ VI-A Common Settings
The experiments used a Universal Robots UR5e manipulator to control rocks with a bucket. Rock position was estimated using an Intel RealSense Depth Camera D435 and bucket force was measured with a Robotiq FT 300-S Force Torque Sensor. The robot controlled the fingertip pose at approximately 100 Hz. Policy learning performance was evaluated in simulation using an Intel Core i9-9900X CPU and GeForce RTX 3090 GPU. All experiments used RL parameters as shown in Table I and the network architecture from [49]. For real-world evaluation, nine types of rocks with different shapes and sizes were used as shown in Table IV. Policy learning employed PPO with parallel environments; $E$ environments are simulated in parallel at all resolutions, which is the maximum number of parallel environments achievable on the current PC setup for the highest resolution environment ($\Delta=10$).
@@
Figure 6: Summaries of analysis: (Left) Performance comparison of PRPD components for (a) different patterns of policy mixture rate, (b) different numbers of grid interval, and (c) different numbers of target middle success rate. The box plots represent the sample numbers at which the policies achieve the target success rate $\hat{\tau}$ for each resolution. The dashed lines denote the total learning time from the initial resolution to the plotted resolution. $\alpha$ denotes the policy mixture rate, which is dynamically updated only in $\alpha=\text{opt.}$.. (Right) (d) Relationship between learning time and number of samples in policy learning required to reach $\hat{\tau}$. Each curve and point of (a), (b), (c), and (d) plots the mean and variance per sample over five experiments (each box plot plots the mean). The learning sample is the value until the target success rate $\hat{\tau}$ is reached at the corresponding resolution (for Ours in (d), it is the summation value until $\hat{\tau}$ is reached at the final resolution $\Delta=10$).
@@ VI-B2 Results
As shown in Fig. 5, PRPD learns control policies in less than one-seventh of the time needed to reach the highest success rate compared to fixed resolution learning. Coarser resolutions result in faster convergence, with a 20-fold difference in learning time between $\Delta=70$ and $\Delta=10$, but they also reduce performance by about 35%. PRPD converges in less than one-fourth of the time compared to simultaneous learning across all resolutions. These results demonstrate that the proposed resolution scheduling framework achieves the fastest policy learning.
@@ VI-C2 Results
VI-D Influence of Different Resolution Scheduling
VI-D1 Settings
VI-D2 Results
VI-E Tradeoff between Time-Efficiency and Sample-Efficiency
VI-E1 Settings
As mentioned in Section II-C, various methods have improved RL’s sample efficiency, but less focus has been placed on “time efficiency.” This section shows that the proposed method outperforms traditional sample-efficient methods in time efficiency. SAC [40] and its ensemble variant E-SAC [31] were used as sample-efficient RL algorithms and tested within the PRPD framework and a fixed-resolution framework with $\Delta=10$.
VI-E2 Results
VI-F Sim-to-Real Policy Transfer
This section validates the sim-to-real policy transfer across diverse real-world rock-excavation environments, including silica sand (low viscosity), kinetic sand (high viscosity), and plastic pellets (blocky shape and large grain size). Table IV shows the evaluation of the transfer policies. The experiment shows that the success rate increases as the resolution improves in the real-world environment for fixed resolution policies, as was demonstrated with the simulation results. PRPD achieves performance comparable to fine-resolution fixed policy learning, indicating no adverse effects from using coarse-resolution simulations and successful sim-to-real transfer. We also evaluated the differences in policy performance during sim-to-real transfer depending on the presence of DR parameters. Table IV shows that policy performance dropped to approximately 50 % without DR. This indicates a reality gap between the simulation and real-world environments, mitigated by applying DR. These results demonstrate that the variable-resolution rock excavation simulator and the proposed policy learning framework achieve approximately 90 % success rates for nine types of real rock environments.
TABLE IV: Sim-to-real experiment of rock excavation: This experiment evaluated nine types of rocks (five differently shaped wooden blocks and three different-sized 3D-printed artificial rocks as shown in rock images) on three types of ground (for each element, three numbers correspond to each environment, silica sand, kinetic sand, and plastic pellets in order). “Scale” indicates approximate rock size. The first-row block shows “Ours” (PRPD) and “w/o DR” (PRPD without DR). The second-row block shows PPO at a fixed resolution. Each value indicates the success number per 20 trials, with the last column showing the average rate. The 20 trials test five policies at four positions (0, 60, 120, and 180 degrees) on a 100-mm radius circumference relative to the bucket’s initial position. Task success is defined as rocks being held in the bucket and the bucket being off the ground.

### v3

@@ I Introduction
[…]
@@
The following are this paper’s main contributions:
  • It proposes a novel policy learning framework, PRPD, which enables time-efficient policy learning by progressively increasing simulation resolution and scheduling tasks from coarse to fine based on policy performance.
  • While the underlying policy optimization and distillation mechanisms are based on existing methods, the novelty of the proposed framework lies in integrating them into a unified structure that enables stable and progressive policy transfer across simulation resolutions. This framework effectively achieves the core novelty of progressive policy distillation across varying simulation resolutions for time-efficient policy learning.
  • The effectiveness of PRPD is demonstrated in a complex particle-based excavation task, achieving a 7-fold reduction in total learning time compared to fixed-resolution training, while maintaining comparable task performance and enabling sim-to-real policy transfer in diverse real-world environments.
@@ II-C Sample Efficient Reinforcement Learning Methods
Several methods have been proposed to improve sample efficiency in reinforcement learning, each targeting different problem settings with distinct assumptions and trade-offs.
Model-based RL: This approach utilizes a learned model of the environment’s dynamics to learn policies without direct environment interaction [26, 27, 28]. While highly sample-efficient in principle, it is susceptible to performance degradation caused by inaccuracies in the learned model, which makes it less reliable in high-dimensional or partially observable domains. Consequently, applying this approach to environments with complex or hard-to-model dynamics remains challenging.
Representation Learning: Representation learning aims to extract informative features from raw observations to improve policy learning efficiency [29, 30, 31]. It enhances sample efficiency by enabling more effective use of each observation, particularly in high-dimensional or partially observable environments. However, it often requires additional training objectives and network modules, which increase computational complexity and may demand task-specific tuning. In a related direction, data augmentation improves sample diversity by transforming observations and has shown notable success in visual RL settings [32, 33, 34]. Nevertheless, its effectiveness depends heavily on the choice of transformation and can be limited in tasks where observations are closely coupled with the underlying dynamics.
Imitation Learning: This approach improves sample efficiency by leveraging expert demonstrations to guide policy learning, thereby reducing the need for extensive exploration and environment interaction [35, 36, 37]. In particular, learning from small demonstration datasets or inferring reward functions from limited supervision can significantly accelerate early-stage training. However, the effectiveness of this approach heavily depends on the quality and diversity of the demonstration data, and it often struggles to generalize beyond the behaviors observed in expert trajectories.
Sample-Efficient RL without External Supervision: Off-policy RL methods, such as Soft Actor-Critic (SAC), improve sample reuse by storing and replaying past interactions from a buffer [38, 39]. Also, ensemble learning improves sample efficiency by leveraging multiple models to enhance model generalization [31, 40, 41]. Although sample-efficient, they typically require large numbers of gradient updates and substantial memory for storing transitions, which may reduce time efficiency. Unlike methods that rely on external supervision, these approaches operate solely on state-action interaction data and require no additional annotation, reward shaping, or auxiliary training objectives.
@@
Various methods have improved RL’s “sample efficiency” by extracting more information from each sample or increasing update times to maximize sample utilization. However, these approaches often come with computational overhead, such as increased calculation time and resource demands by adding learning components, which can compromise “time efficiency,” the overall training time for policy learning.
II-D Comparison with Curriculum Reinforcement Learning
Curriculum Reinforcement Learning (CRL) is an RL method that starts with simple tasks and gradually increases task difficulty [42]. This approach builds on the initial learning experiences from simple tasks and gradually adapts the policies for more complex tasks. Eventually, this progressive increase in difficulty results in a reduction in the number of samples required compared to learning solely from complex tasks [43].
In previous CRL approaches, the learning process is typically organized using a curriculum based on task difficulty, while the simulator resolution remains fixed. For example, some methods restrict the range of goals or initial states during early training and gradually expand this range as the policy stabilizes, thereby guiding the agent toward a policy with better generalization capabilities [44, 45]. Other approaches begin training in simplified environments where external disturbances or noise are intentionally removed, and incrementally introduce real-world complexity to adjust the difficulty of learning [46, 47]. In addition, reward shaping has proven effective as a curriculum design strategy, where the reward function is gradually modified. It typically starts with easily obtainable rewards to promote exploration during early training and then transitions to the original reward structure [48, 49]. Thus, previous CRL approaches primarily rely on task difficulty as the basis for designing the learning curriculum.
Our approach can be interpreted as a type of CRL, as it also involves gradually changing the tasks from simple to complex. However, unlike previous CRL approaches, our method aims to reduce the overall learning time required for policy training in simulation, rather than minimizing the number of samples used for training, which is the typical objective in common CRL. To the best of our knowledge, this is the first approach to treat differences in the spatial resolution of complex simulations as distinct tasks within a curriculum framework.
@@ III-A Reinforcement Learning
$\mathcal{L}_{\theta,\phi}^{\pi,V}=$ $\mathbb{E}_{\mathcal{D}_{i}}\left[\min\{\rho A_{k}(s,a),\text{clip}(\rho,1-\epsilon,1+\epsilon)A_{k}(s,a)\}\right.$ (3)
$\left.-c_{1}(V_{k}(s)-\hat{R})^{2}+c_{2}\pi_{k}(a|s)\log\pi_{k}(a|s)\right],$
@@ III-B Conservative Policy Update in Reinforcement Learning
IV Progressive-Resolution Policy Distillation
Simulator Assumption: To enable policy transfer, our framework follows some simulator assumptions. Policies are designed consistently across different resolutions, even though different simulators typically vary in observations and actions. We assume that resolution-affected elements do not impact observations or actions. The policies’ objectives remain the same, with only the resolution differing, so the reward function is kept identical. This framework supports policy learning in high-resolution simulations, assuming that behavior is less accurate at coarser resolutions and becomes finer as resolution improves. Each resolution level corresponds to a distinct simulation environment configured with different resolution parameters, such as soil particle size. Learning samples are collected independently through interactions between policies and their corresponding environments. We assume that resolution-specific environments can be prepared in advance.
IV-A Resolution Scheduling
IV-A1 Execution Flow
IV-A2 Scheduling Setups
@@
$c_{3}$ Weight of distillation loss 0.5
$c_{4}$ Weight of Q-value loss 1
@@ IV-B2 Policy Update with Conservative Policy Transfer
$\alpha_{k+1}=\alpha_{0}\ \mathbb{E}_{s\sim\mathcal{D}_{i}^{(n)}}$ $\big{[}\mathbb{E}_{a^{\prime}\sim\pi^{(n-1)}(s)}[Q_{k}^{(n)}(s,a^{\prime})]$ (7)
$-\mathbb{E}_{a\sim{\pi}_{k}^{(n)}(s)}[Q_{k}^{(n)}(s,{a})]\big{]},$
where $\alpha_{0}$ is a scaling coefficient, $\mathcal{D}^{(n)}$ is the rollout buffer of the current resolution, and $k$ is an update number. Estimated $\alpha_{k+1}$ is utilized by clipping to $[0,1]$. This weighting scheme increases $\alpha$ when the current policy outperforms the previous policy, thereby making the framework robust to the poor teacher effect [54, 55] by applying distillation only when it is expected to improve the student policy, as suggested in previous work [56]. Finally, the loss function of conservative policy update is defined:
$\mathcal{L}_{\theta^{(n)}}^{\pi^{(n)}}=$ $\mathbb{E}_{(s,a)\sim\mathcal{D}_{i}^{(n)}}\big{[}\text{KL}\{\pi^{(n)}_{k}(s,a)||$ (8)
$(1-\alpha_{k+1})\pi^{(n)}_{k}(s,a)+\alpha_{k+1}\pi^{(n-1)}(s,a)\}\big{]},$
where $\theta^{(n)}$ represents the policy network parameters for the current resolution, and KL denotes the Kullback–Leibler divergence. To estimate the policy mixture rate $\alpha$, PRPD learns an auxiliary value function $Q^{(n)}$, parameterized by $\psi^{(n)}$, which is not used for PPO updates but exclusively for estimating $\alpha$ across resolutions. The value function $Q^{(n)}(s,a)$ is trained independently from the actor-critic structure via a temporal-difference-based loss defined as:
$\mathcal{L}_{\psi^{(n)}}^{Q^{(n)}}=$ $\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}_{i}^{(n)}}\big{[}\{Q^{(n)}_{k}(s,a)-$ (9)
$r+\gamma\mathbb{E}_{a^{\prime}\sim\pi^{(n)}(s^{\prime})}[Q^{(n)}_{k}(s^{\prime},a^{\prime})]\}^{2}\big{]}.$
@@
$\mathcal{L}_{\theta^{(n)},\phi^{(n)},\psi^{(n)}}=\mathcal{L}_{\theta^{(n)},\phi^{(n)}}^{\pi^{(n)},V^{(n)}}+c_{3}\mathcal{L}_{\theta^{(n)}}^{\pi^{(n)}}+c_{4}\mathcal{L}_{\psi^{(n)}}^{Q^{(n)}},$ (10)
@@ V-D Variable Resolution Simulator
TABLE IV: Task Success Rate Comparison at Trained Resolution and at the Finest Resolution ($\Delta=10$)
Trained Resolution $\Delta$ [mm] 70 60 50 40 30 20 10
Trained Resolution [%] $99.4\pm 2.6$ $99.1\pm 3.1$ $98.9\pm 2.3$ $98.3\pm 1.9$ $97.5\pm 5.6$ $96.9\pm 7.2$ $95.8\pm 6.4$
$\Delta=10$ [%] $61.1\pm 8.7$ $68.0\pm 8.3$ $76.0\pm 9.7$ $81.3\pm 5.8$ $85.0\pm 7.5$ $91.1\pm 8.9$ $95.8\pm 6.4$
@@ VI Experiments
We conducted experiments to validate the following objectives. PRPD can learn policies in a shorter learning time compared to previous works (Section VI-B). The conservative policy transfer scheme stabilizes learning in PRPD (Section VI-C). Differences in resolution scheduling affect policy learning (Section VI-D). Sensitivity of PRPD to loss weight scaling (Section VI-E). Trade-off between time efficiency and sample efficiency compared to other RL methods (Section VI-F). The policy trained in the simulator can be applied to various real-world environments (Section VI-G).
@@ VI-A Common Settings
The experiments used a Universal Robots UR5e manipulator to control rocks with a bucket. Rock position was estimated using an Intel RealSense Depth Camera D435 and bucket force was measured with a Robotiq FT 300-S Force Torque Sensor. The robot controlled the fingertip pose at approximately 100 Hz. Policy learning performance was evaluated in simulation using an Intel Core i9-9900X CPU and GeForce RTX 3090 GPU. All experiments used RL parameters as shown in Table I and the network architecture from [57]. For real-world evaluation, nine types of rocks with different shapes and sizes were used as shown in Table V. Policy learning employed PPO with parallel environments; $E$ environments are simulated in parallel at all resolutions, which is the maximum number of parallel environments achievable on the current PC setup for the highest resolution environment ($\Delta=10$). Advantage $A$ was estimated using the generalized advantage estimation scheme [58], and used in the PPO update described in Eq. (3).
@@
Figure 6: Summaries of analysis: (Left) Performance comparison of PRPD components for (a) different patterns of policy mixture rate, (b) different numbers of grid interval, and (c) different numbers of target middle success rate. The box plots represent the sample numbers at which the policies achieve the target success rate $\hat{\tau}$ for each resolution. The dashed lines denote the total learning time from the initial resolution to the plotted resolution. $\alpha$ denotes the policy mixture rate, which is dynamically updated only in $\alpha=\text{opt.}$.. (Right) (d) different scaling parameters of loss function in policy learning required to reach $\hat{\tau}$. Each curve and point of (a), (b), (c), and (d) plots the mean and variance per sample over five experiments (each box plot plots the mean). The learning sample is the value until the target success rate $\hat{\tau}$ is reached at the corresponding resolution (for Ours in (d), it is the summation value until $\hat{\tau}$ is reached at the final resolution $\Delta=10$). * mean $p<0.05$.
Figure 7: Relationship between learning time and number of samples in policy learning: Each point plots the mean and variance per sample over five experiments. Sample number is the summation value until $\hat{\tau}$ is reached at the final resolution $\Delta=10$.
@@ VI-B2 Results
[…]
@@ VI-C2 Results
TABLE V: Sim-to-real experiment of rock excavation: This experiment evaluated nine types of rocks (five differently shaped wooden blocks and three different-sized 3D-printed artificial rocks as shown in rock images) on three types of ground (for each element, three numbers correspond to each environment, silica sand, kinetic sand, and plastic pellets in order). “Scale” indicates approximate rock size. The first-row block shows “Ours” (PRPD) and “w/o DR” (PRPD without DR). The second-row block shows PPO at a fixed resolution. Each value indicates the success number per 20 trials, with the last column showing the average rate. The 20 trials test five policies at four positions (0, 60, 120, and 180 degrees) on a 100-mm radius circumference relative to the bucket’s initial position. Task success is defined as rocks being held in the bucket and the bucket being off the ground.
@@
VI-D Influence of Different Resolution Scheduling
VI-D1 Settings
VI-D2 Results
VI-E Impact of Loss Term Balancing in PRPD
VI-E1 Settings
To evaluate the robustness of the proposed PRPD framework with respect to loss weighting, we conduct a sensitivity analysis on the scale of each loss term in the overall objective as shown in Eq. (10). In this equation, $c_{3}$ and $c_{4}$ control the relative importance of the distillation loss and Q-value loss, respectively. This experiment systematically varies $c_{3}$ and $c_{4}$ to examine how loss balance affects training stability and performance.
VI-E2 Results
As shown in Fig. 6, the proposed framework achieves both sample- and time-efficient learning across a broad range of loss weight settings. Sample efficiency is highest when $c_{3}$ lies between 0.1 and 1, suggesting that moderate incorporation of prior policy guidance is beneficial, whereas excessive emphasis may hinder learning. For $c_{4}$, optimal performance is observed when the value ranges from 0.5 to 1, indicating that $Q$-value estimation supports conservative transfer, though overly large weights may destabilize training. While the framework remains robust to moderate variations, setting $c_{3}$ or $c_{4}$ far outside these ranges can increase the required number of samples by approximately 1.5 times.
VI-F Tradeoff between Time-Efficiency and Sample-Efficiency
VI-F1 Settings
As mentioned in Section II-C, various methods have improved RL’s sample efficiency, but less focus has been placed on “time efficiency.” This section shows that the proposed method outperforms traditional sample-efficient methods in time efficiency. We evaluate Soft Actor-Critic (SAC) [38] and its ensemble-based extension E-SAC [31], along with three recent off-policy algorithms recognized for their sample efficiency and strong empirical performance: Implicit Q-Learning (IQL) [59], Randomized Ensembled Double Q-learning (REDQ) [60], and Truncated Quantile Critics (TQC) [61]. These methods are tested within both the proposed PRPD framework and a fixed-resolution setting with $\Delta=10$.
VI-F2 Results
VI-G Sim-to-Real Policy Transfer
[…]